# Instruct-IPT: All-in-One Image Processing Transformer via Weight Modulation

### v1

@@ Abstract
Due to the unaffordable size and intensive computation costs of low-level vision models, All-in-One models that are designed to address a handful of low-level vision tasks simultaneously have been popular. However, existing All-in-One models are limited in terms of the range of tasks and performance. To overcome these limitations, we propose Instruct-IPT – an All-in-One Image Processing Transformer that could effectively address manifold image restoration tasks with large inter-task gaps, such as denoising, deblurring, deraining, dehazing, and desnowing. Rather than popular feature adaptation methods, we propose weight modulation that adapts weights to specific tasks. Firstly, we figure out task-sensitive weights via a toy experiment and introduce task-specific biases on top of them. Secondly, we conduct rank analysis for a good compression strategy and perform low-rank decomposition on the biases. Thirdly, we propose synchronous training that updates the task-general backbone model and the task-specific biases simultaneously. In this way, the model is instructed to learn general and task-specific knowledge. Via our simple yet effective method that instructs the IPT to be task experts, Instruct-IPT could better cooperate between tasks with distinct characteristics at humble costs. Further, we propose to maneuver Instruct-IPT with text instructions for better user interfaces. We have conducted experiments on Instruct-IPT to demonstrate the effectiveness of our method on manifold tasks, and we have effectively extended our method to diffusion denoisers as well. The code is available at https://github.com/huawei-noah/Pretrained-IPT.
@@ 2 Related Work
Transformer-based Image Restoration Methods. Transformers have pushed image restoration performance further beyond canonical CNN counterparts. Several transformer-based approaches [4, 26, 27, 38, 56] have been proposed to tackle various degradation restoration tasks. Among these methods, IPT [4] first introduced a standard transformer-based framework with multi-heads and multi-tails structure to handle multiple image restoration tasks. SwinIR [31] employed Swin-transformer blocks with a shifted window scheme tailored for restoration tasks, while Uformer [56] combined U-Net and transformer with window-based self-attention to improve model performances. Other transformer-based methods, such as Restormer [60] and IPTv2 [53], leveraged channel self-attention to capture longer-range dependencies.
@@ 3 Method
We have selected a pool of the latest multi-low-level-task adaptation methods as follows. PromptIR [39]: task-specific prompts are injected at each U-Net stage transition; InstructIR [10]: a series of task-specific weights perform channel-wise affine feature mapping; External-Control: inspired by ControlNet [63], a degradation-sensitive parallel encoder network is added that intervenes features in the original backbone. Notably, these adaptation methods all modify intermediate features for certain tasks. We align these feature-based methods to the IPT-V2 [53] model. As an improved version of IPT [4], IPT-V2 [53] achieves outstanding performance on low-level image restoration tasks, and thus being a good base model for our experiment.
@@
PromptIR [39] 34.37 35.13 32.85 30.81
InstructIR [10] 34.31 34.93 33.11 31.06
External-Control 34.38 35.14 32.64 30.57
@@
Table 1: Experimenting feature adaptation methods on distinct tasks. We mix two tasks for training as the baseline, and compare a series of feature adaptation methods against it. Surprisingly, these methods hardly work; they usually sacrifice one task for the other.
@@ 3.2 Efficient Weight Adaptation
Which Weights are Task-Sensitive? Previously, there are some methods that modifies weights for specific tasks [38, 67]. However, these methods universally apply weight modification on all types of weights without evaluating their individual contribution to different tasks.
@@
Layers Cos. Sim.
Img. Processing
Img. Embedder 0.97
Output Layer 0.68
U-Net Layers
Up & DownSampling 0.52
Channel Reduction 0.73
Transformer Layers
QKV Projection 0.72
Post-Attn Projection 0.80
FFN Projection 0.73
LayerNorm 0.99
@@
The Proposed Weight Modulation. The simple idea behind weight modulation is to add task-specific biases on top of IPT weights to adapt them to various tasks. For instance, for a weight $W^{\prime}$ that performs linear mapping on a certain task, we decompose it into the matrix addition of the weight for general restoration task knowledge $W_{general}$ and task-specific modulation bias $B_{task}$ as follows:
$Y=W^{\prime}X=(W_{general}+B_{task})X,$ (1)
@@
$B_{1}\in\mathbb{R}^{n_{\alpha}\times n_{\gamma}}\land B_{2}\in\mathbb{R}^{n_{\gamma}\times n_{\beta}}\quad\text{s.t.}\quad B_{task}=B_{1}B_{2}.$ (2)
@@
$E=\sum^{i=1}_{r}\sigma_{i}^{2}\quad\text{s.t.}\quad\frac{W}{\|W\|_{F}}=\sum^{i=1}_{k}u_{i}^{T}\sigma_{i}v_{i},\quad U^{T}U=I,\quad V^{T}V=I.$ (3)
@@
As a simple yet effective measure, we unfreeze the backbone and train both the backbone and the task-specific biases synchronously. We conduct synchronous training on IPT-V2 and compare it with the conventional two-stage method. Experiment results in Tab. 3 prove the effectiveness of our method.
@@
Table 3: Comparing Synchronous Training with conventional Two-Stage Training (Training & Finetuning). The results demonstrate that Synchronous Training, i.e. training both weight and bias simultaneously, performs better than the Two-Stage process.
Text Commands. Further beyond weight modulation, we inject natural language as commands into the model. In reality, images are usually judged by humans in a subjective manner: the demands of humans might vary under different circumstances. Thus we hope language commands, as a better user interface, could enable complicated demands in an interactive manner. Thanks to the development of Large Language Models (LLMs), we are able to leverage their strong capabilities to generate authentic human commands to simulate the interactive process. The procedures are detailed in Sec. 4.
Figure 3: Qualitative comparisons between Instruct-IPT and PromptIR. We compare the two methods on three tasks: denoising ($\sigma=50$), deraining, and denoising. Our Instruct-IPT could outcompete PromptIR in terms of visual quality.
4 Experiments
Implementation details: Our training set is a large combination of various datasets from multiple tasks, mainly following the datasets in All-in-One image restoration methods [39, 38, 4]. Our training hyperparameters largely follow IPTV2 [53]. We classify all tasks into two parts: elementary and downstream tasks. We hold that denoising and deblurring are elementary tasks that contribute to other downstream tasks. Hence, we perform synchronous training on these two tasks to learn weights for general knowledge and biases for specific knowledge. Accordingly, the training is lengthened by a factor of two. We simply finetune biases for other tasks. We use 8 NVIDIA A100 GPUs.
@@ 4 Experiments
4.1 Comparison with All-in-One Methods
@@
Table 5: Comparing Instruct-IPT with deraining experts. PSNR (on Y) on the Rain100L [59] is reported. The best and second best results are bolded and underlined.
@@
Table 6: Comparing Instruct-IPT with dehazing experts. PSNR on the SOTS Outdoor dataset [25] is reported. The best and second best results are bolded and underlined.
@@
Table 7: Comparing Instruct-IPT with desnowing experts. PSNR on the CSD [7] desnow dataset is reported. The best and second best results are bolded and underlined.
@@ 4.2 Comparison with Single-Task Experts
Tab. 5 to 7 presents a series of comparisons of our proposed Instruct-IPT, trained in an All-in-One manner, against a suite of expert models specifically trained for each individual task. Our results demonstrate that Instruct-IPT achieves competitive performance, approaching the state-of-the-art results of expert models. While not yet surpassing the current best performance, Instruct-IPT ranks among the top-tier techniques, underscoring its efficacy and potential for further development. This analysis highlights the viability of our unified training approach as a compelling alternative to task-specific models, particularly when considering the balance between model generalizability and task-specific performance. Due to page limits, some tables are presented in the Appendix.
@@ 4.3 Extrapolation to Diffusion Models
We evaluate the quality of generated images using four metrics: LPIPS(Learned Perceptual Image Patch Similarity) [64], DISTS(Deep Image Structure and Texture Similarity) [15], FID(Fréchet Inception Distance) [19], and IS(Inception Score) [48]. These metrics assess different aspects of image similarity and quality, offering a comprehensive assessment of image generation quality, and highlighting the perceptual fidelity, structure consistency, distributional similarity, and diversity of generated images. As shown in Tab. 8, we compared the performance of our proposed method against several baseline approaches: inference with original weights, direct fine-tuning, and fine-tuning only task-specific weight biases. Our results demonstrate that our method consistently outperforms these baselines across all evaluated metrics. Fig. 4 presents a qualitative comparison, showcasing our method’s ability to produce more visually compelling and appealing results compared to other baselines.
Inpainting Outpainting
[…]
@@ 5 Conclusion
Existing All-in-One image restoration methods are limited both in terms of task scope and performance. In this paper, we propose an All-in-One method that has outstanding performance on a wide range of image restoration tasks. We start by experimenting with existing feature adaptation methods, which turn out to be ineffective on tasks that are highly different. Then we impose simple task-specific weight modulation on model weights that are sensitive to specific tasks. Due to practicability concerns, we perform rank analysis and figure out a suitable low-rank decomposition strategy for the task-specific bias on weights. Further beyond, we propose synchronous training that enables the model to learn general knowledge in the backbone weights and inject task-specific knowledge to the biases in an automatic fashion. Lastly, we introduce text commands as a good user interface. We introduce the method to IPT and develop Instruct-IPT, which is the SOTA All-in-One image restoration model. The proposed model could achieve supreme performance on various restoration tasks, including denoising, deblurring, deraining, dehazing, and desnowing. We also extended our All-in-One method to diffusion models to demonstrate the generalizability of our method.
Limitations. Due to limited time and computation resources, we have not generalized our method to SOTA Diffusion Transformers.
Broader Impacts & Safeguards. This work contains image generation, which might lead to potential misuse. Though we are not technologically addressing it, we condemn any deliberate misuse of the model by humans.
Appendix

### v2

@@ Abstract
Due to the unaffordable size and intensive computation costs of low-level vision models, All-in-One models that are designed to address a handful of low-level vision tasks simultaneously have been popular. However, existing All-in-One models are limited in terms of the range of tasks and performance. To overcome these limitations, we propose Instruct-IPT – an All-in-One Image Processing Transformer (IPT) that could effectively address manifold image restoration tasks with large inter-task gaps, such as denoising, deblurring, deraining, dehazing, and desnowing. While most research propose feature adaptation methods, we reveal their failure in addressing highly distinct tasks, and suggest weight modulation that adapts weights to specific tasks. Firstly, we search for task-sensitive weights and introduce task-specific biases on top of them. Secondly, we conduct rank analysis for a good compression strategy and perform low-rank decomposition on the biases. Thirdly, we propose synchronous training that updates the task-general backbone model and the task-specific biases simultaneously. In this way, the model is instructed to learn both general and task-specific knowledge. Via our simple yet effective method that instructs the IPT to be task experts, Instruct-IPT could better cooperate between tasks with distinct characteristics at humble costs. As an additional feature, we enable Instruct-IPT to receive human prompts. We have conducted experiments on Instruct-IPT to demonstrate the effectiveness of our method on manifold tasks, and we have effectively extended our method to diffusion denoisers as well. The code is available at https://github.com/huawei-noah/Pretrained-IPT.
@@ 2 Related Work
Transformer-based Image Restoration Methods. Transformers have pushed image restoration performance further beyond canonical CNN counterparts. Several transformer-based approaches [4, 26, 27, 38, 56] have been proposed to tackle various degradation restoration tasks. Among these methods, IPT [4] first introduced a standard transformer-based framework with multi-heads and multi-tails structure to handle multiple image restoration tasks. SwinIR [31] employed Swin-transformer blocks with a shifted window scheme tailored for restoration tasks, while Uformer [56] combined U-Net and transformer with window-based self-attention to improve model performances. Other transformer-based methods, such as Restormer [61] and IPT-V2 [53], leveraged channel self-attention to capture longer-range dependencies.
@@ 3 Method
We have selected a pool of the latest multi-low-level-task adaptation methods as follows. Prompting from PromptIR [39]: task-specific prompts are injected at each U-Net stage transition; Instruct-Channel from InstructIR [10]: a series of task-specific weights perform channel-wise affine feature mapping; External-Control from ControlNet [63]: a degradation-sensitive parallel encoder network is added that intervenes features in the original backbone. Notably, these adaptation methods all modify intermediate features for certain tasks. We align these feature-based methods to the IPT-V2 [53] model. As an improved version of IPT [4], IPT-V2 [53] achieves outstanding performance on low-level image restoration tasks, and thus being a good base model for our experiment.
Layers Similarity
Image Processing
   Image Embedder 0.97
   Output Layer 0.68
U-Net Layers
   UpSampling & DownSampling 0.52
   Channel Reduction 0.73
Transformer Layers
   Query-Key-Value Projection 0.72
   Post-Attention Projection 0.80
   Feed-Forward-Network Projection 0.73
   LayerNorm 0.99
@@
Prompting from PromptIR [39] 34.37 35.13 32.85 30.81
Instruct-Channel from InstructIR [10] 34.31 34.93 33.11 31.06
External-Control from ControlNet [63] 34.38 35.14 32.64 30.57
@@
Table 2: Experimenting feature adaptation methods on distinct tasks. We mix two tasks for training as the baseline, and compare a series of recent feature adaptation methods against it. For fair comparison, all methods are applied to the same backbone and trained for the same # of iterations.
@@ 3.2 Efficient Weight Adaptation
Picking Task-Sensitive Weights. Previously, there are some methods that modifies weights for specific tasks [38, 67]. However, these methods universally apply weight modification on all types of weights without evaluating their individual contribution to different tasks.
@@
The Proposed Weight Modulation. The simple idea behind weight modulation is to add task-specific biases on top of IPT weights to adapt them to various tasks. For instance, for a weight $W^{\prime}$ that performs linear mapping on a certain task, we decompose it into the matrix addition of the weight for general restoration task knowledge $W_{\text{general}}$ and task-specific modulation bias $B_{\text{task}}$ as follows:
$Y=W^{\prime}X=(W_{\text{general}}+B_{\text{task}})X,$ (1)
@@
$B_{1}\in\mathbb{R}^{n_{\alpha}\times n_{\gamma}}\land B_{2}\in\mathbb{R}^{n_{\gamma}\times n_{\beta}}\quad\text{s.t.}\quad B_{\text{task}}=B_{1}B_{2}.$ (2)
@@
[…]
where $\quad U^{T}U=I,V^{T}V=I.$
@@
Table 3: Comparing Synchronous Training with conventional Two-Stage Training (Training & Finetuning). The overall training iterations is the same for all baselines. The results demonstrate that Synchronous Training, i.e. training both weight and bias simultaneously, performs better than the Two-Stage process.
As a simple yet effective measure, we unfreeze the backbone and train both the backbone and the task-specific biases synchronously. We conduct synchronous training on IPT-V2 [53] and compare it with the conventional two-stage method. Experiment results in Tab. 3 prove the effectiveness of our method.
@@
4 Experiments
In this section, we conduct experiments to demonstrate the outstanding performance of the proposed Instruct-IPT model.
Implementation details: Our training set is a large combination of various datasets from multiple tasks, mainly following the datasets in All-in-One image restoration methods [39, 38, 4]. Our training hyperparameters largely follow IPT-V2 [53]. We classify all tasks into two parts: elementary and downstream tasks. We hold that denoising and deblurring are elementary tasks that contribute to other downstream tasks. Hence, we perform synchronous training on these two tasks to learn weights for general knowledge and biases for specific knowledge. Accordingly, the training is lengthened by a factor of two. We simply finetune biases for other tasks. We use 8 NVIDIA A100 GPUs.
Model FLOPs Latency Params
AirNet 19.47G 0.0373s 8.93M
InstructIR 1.02G 0.0259s 15.84M
PromptIR 10.83G 0.0751s 35.59M
Ins.IPT-Tiny (Ours) 0.95G 0.0193s 2.60M
Ins.IPT (Ours) 10.01G 0.0598s 26.49M
Table 5: FLOPs, latency, and # of parameter statistics of baselines. In accordance to baselines, we design two versions of Instruct-IPT to match their sizes.
@@ 4 Experiments
Table 6: Comparing Instruct-IPT with downstream task experts. Upper: Deraining comparison. PSNR (on Y) on the Rain100L [59] is reported. Middle: Dehazing comparison. PSNR on the SOTS Outdoor dataset [25] is reported. Lower: Desnowing comparison. PSNR on the CSD [7] desnow dataset is reported. The best and second best results are bolded and underlined.
Overhead Comparison. We report the FLOPs, latency, and parameter information of mainstream baselines in Tab. 5. As all-in-one works seldom report overhead statistics, most statistics are measured by us from their opensourced codes. To match the models, we propose two versions of Instruct-IPT: the tiny version matches the powerful baseline of InstructIR [10], while the normal version is aligned to PromptIR [39].
4.1 Comparison with All-in-One Methods
Figure 3: A demo of Instruct-IPT instructed by human language. Our method could achieve good image restoration results on various tasks while responding to human language.
@@ 4.2 Comparison with Single-Task Experts
Tab. 6 presents a series of comparisons of our proposed Instruct-IPT, trained in an All-in-One manner, against a suite of expert models specifically trained for each individual task. Our results demonstrate that Instruct-IPT achieves competitive performance, approaching the state-of-the-art results of expert models. While not yet surpassing the current best performance, Instruct-IPT ranks among the top-tier techniques, underscoring its efficacy and potential for further development. This analysis highlights the viability of our unified training approach as a compelling alternative to task-specific models, particularly when considering the balance between model generalizability and task-specific performance. Due to page limits, some tables are presented in the Appendix.
4.3 Human Instruction
Further beyond weight modulation, we inject natural language as commands into the model. In reality, images are usually judged by humans in a subjective manner: the demands of humans might vary under different circumstances. Thus we hope language commands, as a better user interface, could enable complicated demands in an interactive manner. Thanks to the development of Large Language Models (LLMs), we could leverage their strong capabilities to generate a large number of authentic human commands. Then we are able to finetune the additional text encoder with these commands, which responds well to human language. As shown in Fig.  3, the proposed model could respond to various casual instructions from humans. Sometimes, human language can be fairly casual, but Instruct-IPT is still able to respond accurately based on human instructions. This feature enables real-world applications in our daily life.
Figure 4: Qualitative comparisons between Instruct-IPT and competitive baselines. We compare the two methods on three tasks: denoising ($\sigma=50$), deraining, and denoising. Our Instruct-IPT could outcompete baselines by large margins in terms of visual quality.
@@ 4.4 Extrapolation to Diffusion Models
We compared the performance of our proposed method against several baseline approaches: inference with original weights, direct fine-tuning, and fine-tuning only task-specific weight biases. We adopt metrics including LPIPS (Learned Perceptual Image Patch Similarity) [64], DISTS (Deep Image Structure and Texture Similarity) [15], FID (Fréchet Inception Distance) [19], and IS (Inception Score) [48]. Our results demonstrate that our method consistently outperforms these baselines across all evaluated metrics. Due to page limits, we put the table in Appendix. We also present a visual qualitative comparison in the appendix, showcasing our method’s ability to produce more visually compelling and appealing results compared to other baselines.
@@ 5 Conclusion
Existing All-in-One image restoration methods are limited both in terms of task scope and performance. In this paper, we propose an All-in-One method that has outstanding performance on a wide range of image restoration tasks. We start by benchmarking existing task adaptatio methods. We experiment with existing feature adaptation methods on a fair setting, which turn out to be ineffective on tasks that are highly different. Then we impose simple task-specific weight modulation on model weights that are sensitive to specific tasks. Due to practicability concerns, we perform rank analysis and figure out a suitable low-rank decomposition strategy for the task-specific bias on weights. Further beyond, we propose synchronous training that enables the model to learn general knowledge in the backbone weights and inject task-specific knowledge to the biases in an automatic fashion. We introduce the method to IPT and develop Instruct-IPT, which is the SOTA All-in-One image restoration model. The proposed model could achieve supreme performance on various restoration tasks, including denoising, deblurring, deraining, dehazing, and desnowing. Further, we introduce text commands as a good user interface. Beyond conventional regression-based low-level vision tasks, we also extended our All-in-One method to diffusion models to demonstrate the generalizability of our method.
Supplementary Material
@@ Appendix A Additional Experiments
Inpainting Outpainting
[…]
Extrapolation to Diffusion. Due to page limits, we record our contribution of the methods’ extrapolation to diffusion models in this appendix section. We evaluate the quality of generated images using four metrics: LPIPS(Learned Perceptual Image Patch Similarity) [64], DISTS(Deep Image Structure and Texture Similarity) [15], FID(Fréchet Inception Distance) [19], and IS(Inception Score) [48]. These metrics assess different aspects of image similarity and quality, offering a comprehensive assessment of image generation quality, and highlighting the perceptual fidelity, structure consistency, distributional similarity, and diversity of generated images. As shown in Tab. 11, we compared the performance of our proposed method against several baseline approaches: inference with original weights, direct fine-tuning, and fine-tuning only task-specific weight biases. Our results demonstrate that our method consistently outperforms these baselines across all evaluated metrics. We also present a visual qualitative comparison here. Our method could also outcompete baselines in terms of visual quality.
Fig. 5 presents a qualitative comparison, showcasing our method’s ability to produce more visually compelling and appealing results compared to other baselines.